OpenStudy (anonymous):

integral 1/ x[ sqrt (lnx) ]

OpenStudy (anonymous):

What are you asking? Is 1/x ever in sqrt(lnx)? I think that's how satelite interpreted it

OpenStudy (anonymous):

\[\int\limits_{?}^{?} \frac{ 1 }{ x \sqrt{lnx} }dx\]

OpenStudy (anonymous):

you're looking for the limits, to give you what area under the curve?

OpenStudy (anonymous):

NO just a evaluate the integral

OpenStudy (anonymous):

Oh I assumed the ?'s meant find the limits :3